a ) Describe the transformations to the graph of y = x² to obtain y = -2 ( x + 5 )²- 3 .

b ) Graph y = x² . Then apply the transformations in part ( a ) to graph y = -2 ( x + 5 ) ² - 3 .

c ) Determine the domain and range of this transformed function .


1
Expert's answer
2020-12-07T08:00:48-0500

a) Consider the transformed function "y=-2(x+5)^2-3" from the parent function "y=x^2"


i) Multiply "x^2" by 2 to obtain "h(x)=2x^2" in order to stretch the curve vertically as shown in the figure below:





ii) Multiply "h(x)=2x^2" by "-1" to obtain the function "g(x)=-2x^2" in order to reflect the curve about "x" -axis as shown in the figure below:




iii) Shift left by 5 units in order to transformed the function "l(x)=-2(x+5)^2" as shown in the figure below:




iv) Shift vertically down by 3 units to obtain the final transformed function "y=-2(x+5)^2-3" as shown in the figure below:





b) The sketch of the graph in part (a) is as shown in the figure below:





c) Since, the function "y=-2(x+5)^2-3" is defined for all "x" in "R^2" ,


So, the domain of the function is "(-\infin,\infin)"


The graph of the function "y=-2(x+5)^2-3" opens downward, so the maximum value is "-3" .


Therefore, the range of the function is "(-\infin,-3]" .